# Tensorization of neural networks for improved privacy and interpretability

### v2

@@ 2.2.3. Trimming
for all $k\in\{2,\ldots,n-1\}$. For $k=1$ we perform SVD on $\widetilde{\Phi}_{1}(x_{1},\gamma_{1})$ to obtain $\widetilde{\Phi}_{1}(x_{1},\gamma_{1})=B_{1}(x_{1},\alpha_{1})C_{1}(\alpha_{1},\gamma_{1})$, and finally, we set $B_{n}(\beta_{n-1},x_{n}):=\widetilde{\Phi}_{n}(\beta_{n-1},x_{n})$.
@@ 2.3. Continuous case
When $f$ is a continuous function, the main argument remains nearly the same, with just a few necessary adaptations. First, let $f:X_{1}\times\cdots\times X_{n}\to\mathbb{R}$, where $X_{1},\ldots,X_{n}\subset\mathbb{R}$. Typically, $X_{k}=[u,v]$, with $u,v\in\mathbb{R}$, for all $k\in[n]$. It is assumed that $f$ can be approximated by an expansion in a basis of functions formed by the tensor product of one-dimensional functions $\phi_{k}(x)=\{\phi_{k}^{i_{k}}(x)\}_{i_{k}=1}^{d_{k}}$, referred to as the embedding functions. Thus, $f$ is represented as
@@
[…]
@@
To solve Eq. 2.18, if the embedding functions $\{\phi_{k}^{i_{k}}\}_{i_{k}=1}^{d_{k}}$ form an orthonormal basis, tensors $\widecheck{\Phi}_{k}$ can be computed as
[…]
for all $j_{k}\in\{1,\ldots,d_{k}\}$. However, in practice, common embeddings do not form an orthonormal basis, in which case Eq. 2.18 is solved via least-squares. This involves evaluating $\widetilde{\Phi}_{k}$ and $\{\phi_{k}^{i_{k}}\}_{i_{k}=1}^{d_{k}}$ at a set of points $\mathbf{y}_{k}$ to impose the matrix equation
@@ 2.4. Complexity
Let $d=\max_{1\leq k\leq n}d_{k}$. Assuming that the complexity of evaluating $f$ is constant and the cardinality of $\mathbf{y}_{k}$ is a multiple of $d$ for all $k\in\{1,\ldots,n\}$, the Sketching step requires $O(N^{2}dn)$ evaluations. Here, as per the notation used earlier, $N=\max_{1\leq a\leq b\leq n}N_{a:b}$. This complexity includes the SketchForming step since forming sketches is implicitly achieved through the evaluations of $f$ on the selected partitions of pivots. Solving the least-squares problems in Fitting (2.18), by multiplying by the pseudoinverse of the coefficient matrices, can be completed in $O((d+N^{2})d^{2}n)$ time. The SVD in Trimming is performed in $O(N^{3}dn)$ time. Denoting $r=\max_{1\leq k\leq n-1}r_{k}$, the matrix multiplications in SystemForming to construct matrices $A_{k}$ involve $O(N^{2}drn)$ operations. Finally, Solving the CDEs (2.4) requires $O(Nr^{2}(d+1)n)$ time. Altogether, the overall complexity is
@@ 2.5. Relation with TT-CI
In our setting, $\mathbf{x}_{a:b}$ denotes the subset formed by the variables $x_{a}$ to $x_{b}$ of the fixed set of pivots $\mathbf{x}$, and thus the number of unique elements $N_{a:b}$ could vary for each interval. In TT-CI, these partial pivots $\mathbf{x}_{a:b}$, also referred to as interpolation sets, are determined independently by a combinatorial optimization procedure, allowing us to fix $N_{a:b}=N$ for all intervals $a\!:\!b$.
@@
Although we do not formally establish properties that the selected pivots must satisfy to yield good results, our proposal of choosing as pivots points from the training set, combined with the defined left and right sketches, aligns with key results from Refs. [73, 75, 69, 77].
@@
Moreover, Ref. [73] demonstrated that Eq. 2.29 yields a TT representation of $f$ if the interpolation sets $\mathbf{x}_{a:b}$ are nested, i.e., if they satisfy $\mathbf{x}_{1:k}\subset\mathbf{x}_{1:k-1}\times[d_{k}]$ and $\mathbf{x}_{k:n}\subset[d_{k}]\times\mathbf{x}_{k+1:n}$. This condition is inherently satisfied in our approach due to the construction of the interpolation sets from fixed pivots.
The TT-RSS algorithm proposed in this paper, while closely resembling TT-CI, differs fundamentally in its purpose: instead of seeking an accurate approximation of a model $\hat{p}$, TT-RSS aims to construct a TT representation of the underlying distribution $p$. This distinction is essential, as tensorizing intricate models may result in a TT model that, while differing significantly from the original model $\hat{p}$ due to potential mismatches in expressive capacity, still provides a good approximation of $p$. TT-RSS achieves this efficiently through a single sweep, performing SVDs to reveal ranks and employing random projections $U_{k}$ to reduce variance. In contrast, TT-CI aims to accurately approximate $\hat{p}$ through iterative optimization with multiple sweeps to refine interpolation sets, with its rank-revealing adaptation further increasing computational cost via two-site sweeps [69]. Furthermore, when results from TT-RSS are unsatisfactory, rather than iterating with TT-CI’s combinatorial optimization to refine interpolation sets, we can optimize all cores using gradient descent methods, starting from the TT-RSS result, to find better approximations of $p$. This approach motivates a general initialization method for TTs that aligns with the training set distribution and selected embedding, as we detail in Section 5.1.
@@ 3. Performance
(3.1) $\frac{\sum_{i=1}^{\tilde{r}_{k}}s_{i}}{\sum_{j}s_{j}}\geq c,$
where $s_{i}$ are the singular values. The ranks are then determined as $r_{k}=\min(\tilde{r}_{k},r)$.
@@ 3.1.1. Random TT
The first scenario we consider involves a function with an exact TT representation. Specifically, we define functions $f^{n}_{\textsc{tt}}:[d]^{n}\to\mathbb{R}$ as TTs, where $n\in\{100,200,500\}$ represents, as before, the length of the chain, or the number of variables. In all cases, we fix $d=2$ and set a constant rank $\bar{r}=10$. To randomly initialize the cores while avoiding numerical instability (e.g., exploding or vanishing outputs), we construct tensors $G_{k}$ as stacks of orthogonal matrices. For each $x_{k}\in[2]$, the matrix $[G_{k}(\alpha_{k-1},x_{k},\alpha_{k})]_{\alpha_{k-1},\alpha_{k}=1}^{r}$ is orthogonal. Consequently, $[G_{1}(x_{1},\alpha_{1})]_{\alpha_{1}=1}^{r}$ and $[G_{n}(\alpha_{n-1},x_{n})]_{\alpha_{n-1}=1}^{r}$ are unit-norm vectors. As a result, $f^{n}_{\textsc{tt}}(x_{1},\ldots,x_{n})$ represents the scalar product of two unit-norm vectors, ensuring that all outputs lie in the range $[-1,1]$.
Using TT-RSS, we reconstruct an approximation $\hat{f}^{n}_{\textsc{tt}}$ to $f^{n}_{\textsc{tt}}$, varying the number of pivots $N$ from 10 to 20. The pivots $\mathbf{x}$ for the decomposition are selected as random configurations sampled uniformly from the domain $[2]^{n}$. Additionally, we use a different set $\mathbf{s}$ of $M=1000$ test samples to evaluate the reconstruction error. For all experiments, we set $r=N$, as it is assumed that the true rank $\bar{r}$ is unknown and therefore increasing the allowed rank would yield better approximations. Furthermore, we use a very small value for the cumulative percentage parameter, $c=1-10^{-5}$. Since the algorithm’s performance can be influenced by multiple sources of randomness, we repeat the decomposition 10 times for each configuration. The results are shown in the top row of Fig. 3.1.
Figure 3.1. Performance of TT-RSS applied to different functions when varying the number of pivots $N$. From left to right, the columns show: the relative error on the $N$ pivots $\mathbf{x}$ used in TT-RSS, $R(\mathbf{x})$; the relative error on a set of $M$ test samples $\mathbf{s}$ from the functions’ domains, $R(\mathbf{s})$; and the time (in seconds) required to perform the decompositions. For each value of $N$, the decomposition is performed 10 times. The figures display the mean values with error bars at ±$\pm 0.5\sigma$, where $\sigma$ denotes the standard deviation. Additionally, three configurations with different numbers of variables, $n=100$, $n=200$, and $n=500$, are shown in different colors.
In this case, since $f^{n}_{\textsc{tt}}$ already has a TT representation with cores $G_{k}$, the goal is to find a new representation, $\hat{f}^{n}_{\textsc{tt}}$, with cores $\widehat{G}_{k}$, that represent the same TT. As shown in the top-center plot in Fig. 3.1, the relative error approaches zero as the number of pivots increases. Note that the errors displayed in the figure represent the mean values of the 10 decompositions performed. However, in many instances, the errors $R(\mathbf{s})$ are less than $10^{-5}$. In such cases, we verify that the fidelity between $f^{n}_{\textsc{tt}}$ and $\hat{f}^{n}_{\textsc{tt}}$, defined as the scalar product of the normalized tensors,
@@ 3.1.2. Slater Functions
As in the previous example, we consider an input dimension $d=2$. To analyze functions of lengths $n\in\{100,200,500\}$, since $n=ml$, we fix $m=5$ and choose $l\in\{20,40,100\}$, so that we increase the number of discretization steps without changing the number of variables $m$. In all cases, we set $L=10$, with a maximum rank of $r=10$ and $c=1-10^{-5}$. From the uniform distribution over the domain, we select $N\in\{30,40,\ldots,100\}$ pivots, along with $M=1000$ samples for testing.
@@
Although these results are less conclusive than those of the previous case, the bottom row of Fig. 3.1 shows that the errors decrease as $N$ increases. Remarkably, even with a higher discretization level of $l=100$, we achieve test errors on the order of $10^{-2}$ for $N\geq 70$, with computational times ranging between 5 and 10 seconds. As observed earlier, the errors at the pivots are generally smaller than those at other samples. However, since the Slater function does not have an exact TT representation, this behavior is expected to persist unless higher ranks $r$ are allowed, enabling better approximations across all input configurations.
@@
Additionally, it is noteworthy that errors at the pivots tend to increase slightly as $N$ grows. This trend suggests that as the approximation improves for the entire function, the tendency to overfit the pivots diminishes. This observation aligns with the expectation that larger numbers of pivots and increased expressive power allow for more balanced approximations.

### v3

@@ 2.2.3. Trimming
for all $k\in\{2,\ldots,n-1\}$. For $k=1$ we obtain $\widetilde{\Phi}_{1}(x_{1},\gamma_{1})=B_{1}(x_{1},\alpha_{1})C_{1}(\alpha_{1},\gamma_{1})$ by performing SVD on $\widetilde{\Phi}_{1}(x_{1},\gamma_{1})$, and finally, we set $B_{n}(\beta_{n-1},x_{n}):=\widetilde{\Phi}_{n}(\beta_{n-1},x_{n})$.
@@ 2.3. Continuous case
When $f$ is a continuous function, the main argument remains nearly the same, with just a few necessary adaptations. First, let $f:X_{1}\times\cdots\times X_{n}\to\mathbb{R}$, where $X_{1},\ldots,X_{n}\subset\mathbb{R}$. Typically, $X_{k}=[u,v]$, with $u,v\in\mathbb{R}$, for all $k\in[n]$. It is assumed that $f$ can be approximated by an expansion in a basis of functions formed by the tensor product of one-dimensional functions $\phi_{k}(x)=(\phi_{k}^{i_{k}}(x))_{i_{k}\in[d_{k}]}$, referred to as the embedding functions. Thus, $f$ is represented as
@@
(2.19) $\int_{X_{a}\times\cdots\times X_{b}}f(x_{a:b})\delta(x_{a:b}-\mathbf{x}^{i}_{a:b})dx_{a:b}=f(\mathbf{x}^{i}_{a:b}).$
@@
To solve Eq. 2.18, if the embedding functions $\{\phi_{k}^{i_{k}}\}_{i_{k}\in[d_{k}]}$ form an orthonormal basis, tensors $\widecheck{\Phi}_{k}$ can be computed as
(2.22) $\widecheck{\Phi}_{k}(\beta_{k-1},j_{k},\gamma_{k})=\int_{X_{k}}\widetilde{\Phi}_{k}(\beta_{k-1},x_{k},\gamma_{k})\phi_{k}^{j_{k}}(x_{k})dx_{k}$
for all $j_{k}\in[d_{k}]$. However, in practice, common embeddings do not form an orthonormal basis, in which case Eq. 2.18 is solved via least-squares. This involves evaluating $\widetilde{\Phi}_{k}$ and $\{\phi_{k}^{i_{k}}\}_{i_{k}\in[d_{k}]}$ at a set of points $\mathbf{y}_{k}$ to impose the matrix equation
@@ 2.4. Complexity
Let $d=\max_{1\leq k\leq n}d_{k}$. Assuming that the complexity of evaluating $f$ is constant and the cardinality of $\mathbf{y}_{k}$ is a multiple of $d$ for all $k\in[n]$, the Sketching step requires $O(N^{2}dn)$ evaluations. Here, as per the notation used earlier, $N=\max_{1\leq a\leq b\leq n}N_{a:b}$. This complexity includes the SketchBuilding step, since building sketches is implicitly achieved through the evaluations of $f$ on the selected partitions of pivots—namely, on sets of the form $\mathbf{x}_{1:k-1}\times\mathbf{y}_{k}\times\mathbf{x}_{k+1:n}$, which contain $N_{1:k-1}d_{k}N_{k:n}$ elements, for all $k\in[n]$. Solving the least-squares problems in Fitting (2.18), by multiplying by the pseudoinverse of the coefficient matrices, can be completed in $O((d+N^{2})d^{2}n)$ time. The SVD in Trimming is performed in $O(N^{3}dn)$ time. Denoting $r=\max_{1\leq k\leq n-1}r_{k}$, the matrix multiplications in SystemForming to construct matrices $A_{k}$ involve $O(N^{2}drn)$ operations. Finally, Solving the CDEs (2.4) requires $O(Nr^{2}(d+1)n)$ time. Altogether, the overall complexity is
@@ 2.5. Relation with TT-CI
In our setting, $\mathbf{x}_{a:b}$ denotes the subset formed by the variables $x_{a}$ to $x_{b}$ of the fixed set of pivots $\mathbf{x}$, and thus the number of unique elements $N_{a:b}$ could vary for each interval. In TT-CI, these partial pivots $\mathbf{x}_{a:b}$, also referred to as interpolation sets, are determined independently by a combinatorial optimization procedure, allowing us to fix $N_{a:b}=N$ for all intervals a:ba\!:\!b.
@@
Although we do not formally establish properties that the selected pivots must satisfy to yield good results (this is, however, an important question that we leave for future work), our proposal of choosing as pivots points from the training set, combined with the defined left and right sketches, aligns with key results from Refs. [73, 75, 69, 77].
@@
Moreover, Ref. [73] demonstrated that the TT approximation of $f$ obtained from Eq. 2.29 is an exact interpolation of the pivots $\mathbf{x}$ if the interpolation sets $\mathbf{x}_{a:b}$ are nested, i.e., if they satisfy $\mathbf{x}_{1:k}\subset\mathbf{x}_{1:k-1}\times[d_{k}]$ and $\mathbf{x}_{k:n}\subset[d_{k}]\times\mathbf{x}_{k+1:n}$. This condition is inherently satisfied in our approach, as the interpolation sets are constructed from fixed pivots. In contrast, TT-RSS does not yield an exact interpolation, since the approximation is ultimately determined by the SVD in the Trimming subroutine rather than by a cross interpolation decomposition. This is beneficial for our purposes, as we aim to accurately approximate the density $p$, and producing exact values of the model $\hat{p}$ could degrade both the accuracy and privacy of the TT-RSS approximation due to overfitting issues.
Therefore, the TT-RSS algorithm proposed in this paper, while closely resembling TT-CI, differs fundamentally in its purpose: instead of seeking an accurate approximation of a model $\hat{p}$, TT-RSS aims to construct a TT representation of the underlying distribution $p$. This distinction is essential, as tensorizing intricate models may result in a TT model that, while differing significantly from the original model $\hat{p}$ due to potential mismatches in expressive capacity, still provides a good approximation of $p$. TT-RSS achieves this efficiently through a single sweep, performing SVDs to reveal ranks and employing random projections $U_{k}$ to reduce variance. In contrast, TT-CI aims to accurately approximate $\hat{p}$ through iterative optimization with multiple sweeps to refine interpolation sets, with its rank-revealing adaptation further increasing computational cost via two-site sweeps [69].
Furthermore, as previously discussed in Section 2.1, the maxvol step in TT-CI may suffer from ergodicity issues [77], failing to identify good representatives of the function’s support in cases of high sparsity, narrow peaks, or discrete symmetries—situations commonly encountered in high-dimensional machine learning tasks. This limitation renders the TT-CI approach insufficient for the problems we aim to address. On the other hand, even when results from TT-RSS are unsatisfactory, rather than iterating with maxvol to refine interpolation sets, we can optimize all cores using gradient descent methods, starting from the TT-RSS result, to find better approximations of $p$. This approach motivates a general initialization method for TTs that aligns with the training set distribution and selected embedding, as we detail in Section 5.1.
@@ 3. Performance
(3.1) $\frac{\sum_{i=1}^{\tilde{r}_{k}}\sigma_{i}}{\sum_{j}\sigma_{j}}\geq c,$
where $\sigma_{i}$ are the singular values. The ranks are then determined as $r_{k}=\min(\tilde{r}_{k},r)$.
@@ 3.1.1. Random TT
The first scenario we consider involves a function with an exact TT representation. Specifically, we define functions $f^{n}_{\textsc{tt}}:[d]^{n}\to\mathbb{R}$ as TTs, where $n\in\{100,200,500\}$ represents, as before, the length of the chain, or the number of variables. In all cases, we fix $d=2$ and set a constant rank $\bar{r}=10$. To randomly initialize the cores while avoiding numerical instability (e.g., exploding or vanishing outputs), we construct tensors $G_{k}$ as stacks of orthogonal matrices. For each $x_{k}\in[2]$, the matrix $G_{k}(\alpha_{k-1},x_{k},\alpha_{k})]_{\alpha_{k-1},\alpha_{k}\in[\bar{r}]}$ is orthogonal. Consequently, $G_{1}(x_{1},\alpha_{1})]_{\alpha_{1}\in[\bar{r}]}$ and $G_{n}(\alpha_{n-1},x_{n})]_{\alpha_{n-1}\in[\bar{r}]}$ are unit-norm vectors. As a result, $f^{n}_{\textsc{tt}}(x_{1},\ldots,x_{n})$ represents the scalar product of two unit-norm vectors, ensuring that all outputs lie in the range [1,1][-1,1].
Using TT-RSS, we reconstruct an approximation $\hat{f}^{n}_{\textsc{tt}}$ to $f^{n}_{\textsc{tt}}$, varying the number of pivots $N$ from 10 to 35. The pivots $\mathbf{x}$ for the decomposition are selected as random configurations sampled uniformly from the domain $2]^{n}$. Additionally, we use a different set $\mathbf{s}$ of M=1  [^0] M=1\,000 test samples to evaluate the reconstruction error. For all experiments, we set $r=N$, as it is assumed that the true rank $\bar{r}$ is unknown and therefore increasing the allowed rank would yield better approximations. Furthermore, we use a very small value for the cumulative percentage parameter, $c=1-10^{-5}$. Since the algorithm’s performance can be influenced by multiple sources of randomness, we repeat the decomposition 10 times for each configuration. The results are shown in the top row of Fig. 3.1.
In this case, since $f^{n}_{\textsc{tt}}$ already has a TT representation with cores $G_{k}$, the goal is to find a new representation, $\hat{f}^{n}_{\textsc{tt}}$, with cores $\widehat{G}_{k}$, that represent the same TT. As shown in the top-center plot in Fig. 3.1, the relative error approaches zero as the number of pivots increases. Note that the errors displayed in the figure represent the mean values of the 10 decompositions performed. However, in many instances, the errors $R(\mathbf{s})$ are less than $10^{-14}$. In such cases, we verify that the fidelity between $f^{n}_{\textsc{tt}}$ and $\hat{f}^{n}_{\textsc{tt}}$, defined as the scalar product of the normalized tensors,
@@
Figure 3.1. Performance of TT-RSS applied to different functions when varying the number of pivots NN (x-axis). From left to right, the columns show: the relative error on the NN pivots $\mathbf{x}$ used in TT-RSS, $R(\mathbf{x})$; the relative error on a set of $M$ test samples $\mathbf{s}$ from the functions’ domains, $R(\mathbf{s})$; and the time (in seconds) required to perform the decompositions. For each value of $N$, the decomposition is performed 10 times. The figures display the mean values with error bars at $\pm 0.5\sigma$, where $\sigma$ denotes the standard deviation. Additionally, three configurations with different numbers of variables, $n=100$, $n=200$, and $n=500$, are shown in different colors. All computations are performed in double precision.
@@ 3.1.2. Slater Functions
As in the previous example, we consider an input dimension $d=2$. To analyze functions of lengths $n\in\{100,200,500\}$, since $n=ml$, we fix $m=5$ and choose $l\in\{20,40,100\}$, so that we increase the number of discretization steps without changing the number of variables $m$. In all cases, we set $L=10$, with a maximum rank of $r=10$ and $c=1-10^{-5}$. From the uniform distribution over the domain, we select $N\in\{30,40,\ldots,100\}$ pivots, along with M=1  [^0] M=1\,000 samples for testing.
@@
Although these results are less conclusive than those of the previous case, the bottom row of Fig. 3.1 shows that the errors decrease as NN increases. Remarkably, even with a higher discretization level of $l=100$, we achieve test errors on the order of $10^{-1}$ for $N\geq 80$, with computational times ranging between 125 and 250 seconds. As observed earlier, the errors at the pivots are generally smaller than those at other samples. However, since the Slater function does not have an exact TT representation, this behavior is expected to persist unless higher ranks $r$ are allowed, enabling better approximations across all input configurations.
@@
Additionally, it is noteworthy that errors at the pivots tend to increase slightly as NN grows. This trend suggests that as the approximation improves for the entire function, the tendency to overfit the pivots diminishes. This observation aligns with the expectation that larger numbers of pivots and increased expressive power allow for more balanced approximations.
3.1.3. Comparison with TT-CI
Table 3.1. Performance comparison of TT-RSS and TT-CI applied to different functions with varying numbers of variables $n$ and pivots NN. The table reports the relative error on a set of $M$ test samples $\mathbf{s}$ from the functions’ domains, $R(\mathbf{s})$, and the time (in seconds) required to perform the decompositions. For TT-CI, the reported time corresponds to the sweeps, while the value in parentheses includes the total time of the full algorithm, including preliminary steps for decomposition setup. Each decomposition is repeated 10 times, and the median values are shown. All computations are performed in double precision.
  $f$ $n$ Tensorization algorithm Error on test Time in secs.
  $T$-RSS ($N=20$) $3.60\times 10^{-1}$ $1.97\times 10^{0}$
100 $T$-RSS ($N=35$) $4.26\times 10^{-15}$ $4.25\times 10^{0}$
Random TT $T$-CI $7.56\times 10^{-15}$ $3.83\times 10^{1}$ ($4.73\times 10^{2}$)
$T$-RSS ($N=20$) $9.23\times 10^{-15}$ $7.08\times 10^{0}$
200 $T$-RSS ($N=35$) $6.56\times 10^{-15}$ $1.89\times 10^{1}$
$T$-CI $1.08\times 10^{-14}$ $1.37\times 10^{2}$ ($2.53\times 10^{3}$)
$T$-RSS ($N=50$) $8.06\times 10^{-1}$ $7.70\times 10^{0}$
100 $T$-RSS ($N=100$) $1.58\times 10^{-1}$ $2.76\times 10^{1}$
$T$-CI $6.27\times 10^{-1}$ $4.09\times 10^{1}$ ($4.71\times 10^{2}$)
$T$-RSS ($N=50$) $3.69\times 10^{-1}$ $1.76\times 10^{1}$
Slater functions 200 $T$-RSS ($N=100$) $2.12\times 10^{-1}$ $5.57\times 10^{1}$
$T$-CI $5.63\times 10^{-1}$ $1.34\times 10^{2}$ ($2.53\times 10^{3}$)
$T$-RSS ($N=50$) $3.61\times 10^{-1}$ $5.22\times 10^{1}$
500 $T$-RSS ($N=100$) $2.29\times 10^{-1}$ $1.94\times 10^{2}$
$T$-CI $6.79\times 10^{-1}$ $3.30\times 10^{2}$ ($1.32\times 10^{4}$)
For comparison with prior tensorization techniques, Table 3.1 shows the performance of TT-RSS and TT-CI—using the implementation from the tntorch package [111]—applied to the previously described functions, under a reduced set of configurations. Specifically, we tensorize random TTs for $n\in\{100,200\}$ and Slater functions for $l\in\{20,40,100\}$ (equivalently, $n\in\{100,200,500\}$), using two pivot set sizes: $N\in\{20,35\}$ for random TTs, and $N\in\{50,100\}$ for Slater functions. For TT-CI, we perform a single round of sweeps along the TT chain—left-to-right and right-to-left—and, importantly, we specify the TT ranks in advance: $r=\bar{r}$ for random TTs, and $r=10$ for Slater functions. That is, TT-CI is given an advantage to yield results that are more directly comparable.
As noted in previous subsections, the randomness in TT-RSS, mainly due to pivot selection, can cause significant variability in results. For instance, with $N=20$ pivots in random TTs, test errors typically concentrate around $10^{-14}$ when the TT is exactly recovered, or around $10^{-1}$ when it is not. Thus, while Fig. 3.1 is useful to observe general trends in error and time as the number of pivots increases, it does not reflect typical performance. To provide a fairer comparison, Table 3.1 reports median values: specifically, the median test errors and median tensorization times (in seconds). For TT-CI, which exhibits less variability due to its more exhaustive optimization, the medians closely match the means. The time costs shown for TT-CI correspond to the sweeps, with the full algorithm time—including pivot preparation and decomposition setup—given in parentheses. For TT-RSS, we report the total time for the full algorithm, which involves a single left-to-right sweep.
In this setting, TT-RSS achieves lower errors in significantly less time. As already shown in Fig. 3.1, increasing the number of pivots reduces the approximation error, yet in some cases, even with the smallest pivot sets, TT-RSS yields lower errors than TT-CI, with notably better time performance. This is not due to running only a single sweep in TT-CI: across all experiments in this work, we observed that once the number of pivots is fixed, TT-CI typically achieves good approximations in a single sweep, and additional sweeps do not significantly improve accuracy. We thus conclude that our pivot selection strategy is sufficient to produce results comparable to TT-CI’s more exhaustive search, while requiring significantly less time for decomposition—especially when compared to the full runtime of TT-CI (shown in parentheses), which often increases by one or two orders of magnitude, rendering it impractical for higher-dimensional functions.
Additionally, as previously mentioned, TT-CI is given a strong advantage here by assuming prior knowledge of the TT ranks. In practice, ranks are usually unknown and must be determined adaptively, as is done in TT-RSS via the SVD. In contrast, TT-CI typically uses a fixed number of pivots per sweep, optimized via the maxvol routine. If the resulting error is too large, the number of pivots—and hence the rank—is increased in subsequent sweeps to enhance expressivity. A realistic use of TT-CI would thus involve multiple sweeps until the desired number of pivots—and hence, the target accuracy—is achieved, substantially increasing the total runtime. Although recent work has proposed rank-revealing versions of TT-CI [77], TT-RSS remains a more favorable alternative, achieving comparable—or in most cases superior—results in a single, efficient left-to-right sweep.